# Online Estimation with Rolling Validation: Adaptive Nonparametric Estimation with Streaming Data

### v3

@@ 8 Discussion
$\left\|\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-\sum_{i=1}^{n}E[u_{i}]\right\|_{2}^{2}$
$=$ $\left\|\sum_{i=1}^{n}\left(E[u_{i}|F^{i-1}]-E[u_{i}|F^{i-2}]+E[u_{i}|F^{i-2}]-E[u_{i}|F^{i-3}]+\cdot\cdot\cdot+E[u_{i}|F^{1}]-E[u_{i}]\right)\right\|_{2}^{2}$
$=$ $\left\|\sum_{i=1}^{n}\sum_{j=1}^{i-1}E\left[u_{i}\mid F^{j}\right]-E\left[u_{i}\mid F^{j-1}\right]\right\|_{2}^{2}$
$=$ $\left\|\sum_{j=1}^{n-1}\sum_{i=j+1}^{n}E\left[u_{i}\mid F^{j}\right]-E\left[u_{i}\mid F^{j-1}\right]\right\|_{2}^{2}$
=$\stackrel{{\scriptstyle(I)}}{{=}}$ $\sum_{j=1}^{n-1}\left\|\sum_{i=j+1}^{n}E[u_{i}|F^{j}]-E[u_{i}|F^{j-1}]\right\|_{2}^{2}$
=$\stackrel{{\scriptstyle(II)}}{{=}}$ $\sum_{j=1}^{n-1}\left\|\sum_{i=j+1}^{n}E[\nabla_{j}u_{i}|F^{j}]\right\|_{2}^{2}\stackrel{{\scriptstyle(III)}}{{\leq}}\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}\|E[\nabla_{j}u_{i}|F^{j}]\|_{2}\right)^{2}\,.$ (54)
In step $(I)$ we used that $M_{j}=\sum_{i=j+1}^{n}E[u_{i}|F^{j}]-E[u_{i}|F^{j-1}]$ is a martingale-difference sequence w.r.t. to the filtration $F^{j}$. Therefore $E[M_{l}M_{k}]=0$ for all $l\neq k$. In step $(II)$ we used the definition of the difference operator $\nabla_{j}$ and realize that $E(u_{i}(Z_{i})|F^{j-1})=E(u_{i}(Z_{i}^{j})|F^{j})$. Step $(III)$ is triangular inequality.
@@
$E[\nabla_{j}u_{i}|F^{j}]$
$=$ $i^{\xi}E[\{\nabla_{j}\hat{f}_{i-1}(X_{i})\}\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})+\hat{f}_{i-1}^{(j)}(X_{i})-f_{0}(X_{i})\}|F^{j}]$
$\leq$ $i^{\xi}\sqrt{E[\{\nabla_{j}\hat{f}_{i-1}(X_{i})\}^{2}|F^{j}]}\sqrt{E[\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})+\hat{f}_{i-1}^{(j)}(X_{i})-f_{0}(X_{i})\}^{2}|F^{j}]}\,.$
Under 4.5 that $\sqrt{E[\{\nabla_{j}\hat{f}_{i-1}(X_{i})\}^{2}|F^{j}]}$ can be bounded by $\Xi_{i}=D_{2}i^{-b}$ almost surely, we can continue (54) as following:
$\sum_{i=j+1}^{n}\left\|E[\nabla_{j}u_{i}|F^{j}]\right\|_{2}$
$\leq$ $\sum_{i=j+1}^{n}i^{\xi}\left\|\sqrt{E[\{\nabla_{j}\hat{f}_{i-1}(X_{i})\}^{2}|F^{j}]}\sqrt{E[\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})+\hat{f}_{i-1}^{(j)}(X_{i})-f_{0}(X_{i})\}^{2}|F^{j}]}\right\|_{2}$
$\leq$ $\sum_{i=j+1}^{n}i^{\xi}\Xi_{i}\left\|\sqrt{E[\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})+\hat{f}_{i-1}^{(j)}(X_{i})-f_{0}(X_{i})\}^{2}|F^{j}]}\right\|_{2}$
@@
$\sum_{i=j+1}^{n}\left\|E[\nabla_{j}u_{i}|F^{j}]\right\|_{2}\lesssim\sum_{i=j+1}^{n}i^{-c},\text{ where }c:=(a/2+b-\xi)\,.$
@@
    $\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}\left\|E[\nabla_{j}u_{i}|F^{j}]\right\|_{2}\right)^{2}\lesssim\sum_{j=1}^{n-1}(j^{1-c}-n^{1-c})^{2}\lesssim n^{3-2c}\vee\log n$
@@
    $c_{n}^{-1}\left\|\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-\sum_{i=1}^{n}E[u_{i}]\right\|_{2}\lesssim c_{n}^{-1}n^{3/2-c}\vee c_{n}^{-1}\log^{1/2}n$
@@
    $\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}\left\|E[\nabla_{j}u_{i}|F^{j}]\right\|_{2}\right)^{2}$
@@
    $\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}\left\|E[\nabla_{j}u_{i}|F^{j}]\right\|_{2}\right)^{2}\lesssim\sum_{j=1}^{n-1}(n^{1-c}-j^{1-c})^{2}\lesssim n^{3-2c}$
    Similar to case 1, we have $c_{n}^{-1}\left\|\sum_{i=1}^{n}E[u_{i}|F^{i-1}]-\sum_{i=1}^{n}E[u_{i}]\right\|_{2}\rightarrow 0$ as $n\rightarrow\infty$. ∎
@@
Goal is estimating $\mu=E[Y]$, our estimator at step $i$ is $\hat{\mu}_{i}=i^{-1}\sum_{l=1}^{i}Y_{i}$. In this case, $u_{i}=i^{\xi}(\hat{\mu}_{i}-\mu)^{2}$. The random quantity of interest is
$\mathcal{D}_{n}=\sum_{i=1}^{n}E\left[u_{i}\mid F^{i-1}\right]-E\left[u_{i}\right].$
We can explicitly calculate the perturbed one version of $u_{i+1}$:
$\nabla_{j}u_{i+1}$ $=i^{\xi}\left(\hat{\mu}_{i}-\hat{\mu}_{i}^{j}\right)\left(\hat{\mu}_{i}+\hat{\mu}_{i}^{j}-2\mu\right)$ (56)
@@
Take conditional expectation
$E\left[\nabla_{j}u_{i+1}\mid F^{j}\right]=i^{\xi-2}S_{j},$ (57)
@@
$S_{j}$ $=2\left(Y_{j}-\mu\right)\left(\sum_{l=1}^{j-1}Y_{l}-j\mu\right)+\left(Y_{j}^{2}-\alpha^{2}\right)$ (58)
$\alpha^{2}$ $=E[Y^{2}].$
@@
$\sum_{i=j+1}^{n}E\left[\nabla_{j}u_{i}\mid F^{j}\right]=S_{j}\left(\sum_{i=j+1}^{n}(i-1)^{\xi-2}\right)=:S_{j}A(j;n).$ (59)
Treat LHS of Equation 59 as a random variable and we calculate its second moment
$E\left[\left(\sum_{i=j+1}^{n}E\left[\nabla_{j}u_{i}\mid F^{j}\right]\right)^{2}\right]=A^{2}(j;n)E[(S_{j})^{2}].$ (60)
@@
$E[(S_{j})^{2}]$ $=4{\rm var}^{2}(Y)j+(E[Y^{4}]-4\mu E[Y^{3}]-5\alpha^{4}-8\mu^{4}+16\mu^{2}\alpha^{2})$ (61)
@@
$\operatorname{var}\left(\mathcal{D}_{n}\right)$ (62)
$=$ $\sum_{j=1}^{n-1}E\left[\left(\sum_{i=j+1}^{n}E\left[\nabla_{j}u_{i}\mid F^{j}\right]\right)^{2}\right]$
@@
Similarly, we can show the $\sum_{j=1}^{n-1}A(j;n)D$ term is greater than a negative constant. This concludes that $\operatorname{var}\left(\mathcal{D}_{n}\right)$ is greater than a constant times $\log n$ for large enough $n$. This corresponds to the $a=b=1$, $\xi=0$ case. Compared with the bound in Lemma 4.4, we know the order of $c_{n}$ cannot be improved in general: We required $c_{n}^{2}$ to be slightly larger than $\log n$, and the latter matches the lower bound of ${\rm var}(\mathcal{D}_{n})$.
@@
This corresponds to the $a=b=1$, $\xi=2$ case and we can compare the bound in Lemma 4.4.
In Figure D.1, we present the histograms of $\mathcal{D}_{n}$ for the mean estimation example with $\xi=1$ and $2$. The variances can be shown to be larger than $n^{2}$ and $n^{4}$, respectively, using the argument above. After dividing them by their standard deviation (SD)—more accurately, a sequence of the same order as their SD lower bounds—we observe their normalized distributions do not visually vary between different sample sizes (and are close to a tamed continuous distribution). Given the above evidence, we conjecture $c_{n}$ in Lemma 4.4 need to be of order larger than $\sqrt{{\rm var}(\mathcal{D}_{n})}$ to have the probability in (21) converging to zero.
@@
$E\left[\nabla_{j}u_{i+1}\mid F^{j}\right]=i^{\xi-2}S_{j}$ (65)
@@
$S_{j}=2\left(Y_{j}-\mu\right)\left(\sum_{l=1}^{j-1}Y_{l}-j\mu\right)+\left(Y_{j}^{2}-\alpha^{2}\right).$ (66)
@@
$E\left[\nabla_{j}u_{i+1}\mid F^{j}\right]$ (67)
$=i^{\xi-2}E\left[\left(Y_{j}-Y_{j}^{\prime}\right)\left(2\sum_{l=1,l\neq j}^{i}Y_{l}+Y_{j}+Y_{j}^{\prime}-2i\mu\right)\mid F^{j}\right]$
$=i^{\xi-2}Y_{j}E\left[2\sum_{l=1,l\neq j}^{i}Y_{l}+Y_{j}+Y_{j}^{\prime}-2i\mu\mid F^{j}\right]$
$\quad-i^{\xi-2}E\left[Y_{j}^{\prime}\left(2\sum_{l=1,l\neq j}^{i}Y_{l}+Y_{j}+Y_{j}^{\prime}-2i\mu\right)\mid F^{j}\right]$
@@ Appendix E Consistency under Diverging Candidate Cardinality
We define $H_{n}=\sum_{i=1}i^{\xi}\min_{k\in K_{i}}E[r(h_{i-1}^{(k)})]$ as the accumulated error of the best alternative models.
pr( correct selection )$\operatorname{pr}(\text{ correct selection })$ (68)
$=$ $\operatorname{pr}\left(\inf_{k\in K_{n}}\sum_{i=1}^{n}i^{\xi}\left\{h^{(k)}_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\right)$
$\geq$ $\operatorname{pr}\left(\inf_{k}\sum_{i=1}^{n}i^{\xi}\left\{h^{(k)}_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\text{ and }\right.$
@@
$=$ $1-\operatorname{pr}\left(\inf_{k}\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right.\text{ or }$
@@
$\geq$ $1-\operatorname{pr}\left(\inf_{k}\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$
$\quad-\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$
$\geq$ $1-\sum_{k\in[K_{n}]}\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$
$\quad-\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right).$
Define $H_{n}^{(k)}=\sum_{i=1}^{n}i^{\xi}E[r(h_{i-1}^{(k)})]$, for each term in the $\sum_{k\in[K_{n}]}$ summation:
$\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$ (69)
$\leq\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}^{(k)}\right)$
$\leq\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}^{(k)}\right|\geq\delta H_{n}^{(k)}\right)$
$\stackrel{{\scriptstyle(I)}}{{\leq}}\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}^{(k)}\right|\geq\delta\sum_{i=1}^{n}i^{-a_{k}+\xi}M_{k}\right)$
$\leq\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h^{(k)}_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}^{(k)}\right|\geq\delta n^{1-a_{k}+\xi}\tilde{M}_{k}/2\right)$
@@
$\stackrel{{\scriptstyle(II)}}{{\leq}}\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)+$
$\quad\operatorname{pr}\left(\left|\sum_{i=1}^{n}u_{i}^{(k)}-E\left[u_{i}^{(k)}\mid F^{i-1}\right]\right|\geq c_{n}\right)+$
$\quad\operatorname{pr}\left(\left|\sum_{i=1}^{n}E\left[u_{i}^{(k)}\mid F^{i-1}\right]-H_{n}^{(k)}\right|\geq c_{n}\right).$
@@
$\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)\lesssim\exp(-K^{\prime}(\mathcal{C}^{(k)}_{n})^{\theta^{\prime}}),$ (70)
@@
$\mathcal{C}^{(k)}_{n}=n^{(1-a_{k})/2}\wedge\log^{-1}n\left(n^{b_{k}-a_{k}/2-1/2}\wedge n^{1-a_{k}+\xi}\right).$ (71)
@@
$\log^{-1}n\cdot n^{\mathcal{C}_{n}/2}\lesssim\mathcal{C}^{(k)}_{n}.$ (72)

### v4

@@ Acknowledgments
$\left\|\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\sum_{i=1}^{n}\mathbb{E}[u_{i}]\right\|_{2}^{2}$
$=$ $\left\|\sum_{i=1}^{n}\left(\mathbb{E}[u_{i}|F^{i-1}]-\mathbb{E}[u_{i}|F^{i-2}]+\mathbb{E}[u_{i}|F^{i-2}]-\mathbb{E}[u_{i}|F^{i-3}]+\cdot\cdot\cdot+\mathbb{E}[u_{i}|F^{1}]-\mathbb{E}[u_{i}]\right)\right\|_{2}^{2}$
$=$ $\left\|\sum_{i=1}^{n}\sum_{j=1}^{i-1}\mathbb{E}\left[u_{i}\mid F^{j}\right]-\mathbb{E}\left[u_{i}\mid F^{j-1}\right]\right\|_{2}^{2}$
$=$ $\left\|\sum_{j=1}^{n-1}\sum_{i=j+1}^{n}\mathbb{E}\left[u_{i}\mid F^{j}\right]-\mathbb{E}\left[u_{i}\mid F^{j-1}\right]\right\|_{2}^{2}$
=$\stackrel{{\scriptstyle(I)}}{{=}}$ $\sum_{j=1}^{n-1}\left\|\sum_{i=j+1}^{n}\mathbb{E}[u_{i}|F^{j}]-\mathbb{E}[u_{i}|F^{j-1}]\right\|_{2}^{2}$
=$\stackrel{{\scriptstyle(II)}}{{=}}$ $\sum_{j=1}^{n-1}\left\|\sum_{i=j+1}^{n}\mathbb{E}[\nabla_{j}u_{i}|F^{j}]\right\|_{2}^{2}\stackrel{{\scriptstyle(III)}}{{\leq}}\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}\|\mathbb{E}[\nabla_{j}u_{i}|F^{j}]\|_{2}\right)^{2}\,.$ (36)
In step $(I)$ we used that $M_{j}=\sum_{i=j+1}^{n}\mathbb{E}[u_{i}|F^{j}]-\mathbb{E}[u_{i}|F^{j-1}]$ is a martingale-difference sequence w.r.t. to the filtration $F^{j}$. Therefore $\mathbb{E}[M_{l}M_{k}]=0$ for all $l\neq k$. In step $(II)$ we used the definition of the difference operator $\nabla_{j}$ and realize that $E(u_{i}(Z_{i})|F^{j-1})=E(u_{i}(Z_{i}^{j})|F^{j})$. Step $(III)$ is triangular inequality.
@@
$\mathbb{E}[\nabla_{j}u_{i}|F^{j}]$
$=$ $i^{\xi}\mathbb{E}[\{\nabla_{j}\hat{f}_{i-1}(X_{i})\}\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})+\hat{f}_{i-1}^{(j)}(X_{i})-f_{0}(X_{i})\}|F^{j}]$
$\leq$ $i^{\xi}\sqrt{\mathbb{E}[\{\nabla_{j}\hat{f}_{i-1}(X_{i})\}^{2}|F^{j}]}\sqrt{\mathbb{E}[\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})+\hat{f}_{i-1}^{(j)}(X_{i})-f_{0}(X_{i})\}^{2}|F^{j}]}\,.$
Under 4.5 that $\sqrt{\mathbb{E}[\{\nabla_{j}\hat{f}_{i-1}(X_{i})\}^{2}|F^{j}]}$ can be bounded by $\Xi_{i}=D_{2}i^{-b}$ almost surely, we can continue (36) as following:
$\sum_{i=j+1}^{n}\left\|\mathbb{E}[\nabla_{j}u_{i}|F^{j}]\right\|_{2}$
$\leq$ $\sum_{i=j+1}^{n}i^{\xi}\left\|\sqrt{\mathbb{E}[\{\nabla_{j}\hat{f}_{i-1}(X_{i})\}^{2}|F^{j}]}\sqrt{\mathbb{E}[\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})+\hat{f}_{i-1}^{(j)}(X_{i})-f_{0}(X_{i})\}^{2}|F^{j}]}\right\|_{2}$
$\leq$ $\sum_{i=j+1}^{n}i^{\xi}\Xi_{i}\left\|\sqrt{\mathbb{E}[\{\hat{f}_{i-1}(X_{i})-f_{0}(X_{i})+\hat{f}_{i-1}^{(j)}(X_{i})-f_{0}(X_{i})\}^{2}|F^{j}]}\right\|_{2}$
@@
$\sum_{i=j+1}^{n}\left\|\mathbb{E}[\nabla_{j}u_{i}|F^{j}]\right\|_{2}\lesssim\sum_{i=j+1}^{n}i^{-c},\text{ where }c:=(a/2+b-\xi)\,.$
@@
    $\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}\left\|\mathbb{E}[\nabla_{j}u_{i}|F^{j}]\right\|_{2}\right)^{2}\lesssim\sum_{j=1}^{n-1}(j^{1-c}-n^{1-c})^{2}\lesssim n^{3-2c}\vee\log n$
@@
    $c_{n}^{-1}\left\|\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\sum_{i=1}^{n}\mathbb{E}[u_{i}]\right\|_{2}\lesssim c_{n}^{-1}n^{3/2-c}\vee c_{n}^{-1}\log^{1/2}n$
@@
    $\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}\left\|\mathbb{E}[\nabla_{j}u_{i}|F^{j}]\right\|_{2}\right)^{2}$
@@
    $\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}\left\|\mathbb{E}[\nabla_{j}u_{i}|F^{j}]\right\|_{2}\right)^{2}\lesssim\sum_{j=1}^{n-1}(n^{1-c}-j^{1-c})^{2}\lesssim n^{3-2c}$
    Similar to case 1, we have $c_{n}^{-1}\left\|\sum_{i=1}^{n}\mathbb{E}[u_{i}|F^{i-1}]-\sum_{i=1}^{n}\mathbb{E}[u_{i}]\right\|_{2}\rightarrow 0$ as $n\rightarrow\infty$. ∎
@@
Goal is estimating $\mu=\mathbb{E}[Y]$, our estimator at step $i$ is $\hat{\mu}_{i}=i^{-1}\sum_{l=1}^{i}Y_{i}$. In this case, $u_{i}=i^{\xi}(\hat{\mu}_{i}-\mu)^{2}$. The random quantity of interest is
$\mathcal{D}_{n}=\sum_{i=1}^{n}\mathbb{E}\left[u_{i}\mid F^{i-1}\right]-\mathbb{E}\left[u_{i}\right].$
We can explicitly calculate the perturb-one version of $u_{i+1}$:
$\nabla_{j}u_{i+1}$ $=i^{\xi}\left(\hat{\mu}_{i}-\hat{\mu}_{i}^{j}\right)\left(\hat{\mu}_{i}+\hat{\mu}_{i}^{j}-2\mu\right)$
@@
Take the conditional expectation
$\mathbb{E}\left[\nabla_{j}u_{i+1}\mid F^{j}\right]=i^{\xi-2}S_{j},$ (38)
@@
$S_{j}$ $=2\left(Y_{j}-\mu\right)\left(\sum_{l=1}^{j-1}Y_{l}-j\mu\right)+\left(Y_{j}^{2}-\alpha^{2}\right)$
$\alpha^{2}$ $=\mathbb{E}[Y^{2}].$
@@
$\sum_{i=j+1}^{n}\mathbb{E}\left[\nabla_{j}u_{i}\mid F^{j}\right]=S_{j}\left(\sum_{i=j+1}^{n}(i-1)^{\xi-2}\right)=:S_{j}A(j;n).$ (39)
Treat LHS of Equation 39 as a random variable, and we calculate its second moment
$\mathbb{E}\left[\left(\sum_{i=j+1}^{n}\mathbb{E}\left[\nabla_{j}u_{i}\mid F^{j}\right]\right)^{2}\right]=A^{2}(j;n)\mathbb{E}[(S_{j})^{2}].$
@@
$\mathbb{E}[(S_{j})^{2}]$ $=4{\rm var}^{2}(Y)j+(\mathbb{E}[Y^{4}]-4\mu\mathbb{E}[Y^{3}]-5\alpha^{4}-8\mu^{4}+16\mu^{2}\alpha^{2})$
@@
$\operatorname{var}\left(\mathcal{D}_{n}\right)$
$=$ $\sum_{j=1}^{n-1}\mathbb{E}\left[\left(\sum_{i=j+1}^{n}\mathbb{E}\left[\nabla_{j}u_{i}\mid F^{j}\right]\right)^{2}\right]$
@@
Similarly, we can show that the $\sum_{j=1}^{n-1}A(j;n)D$ term is greater than a negative constant. This concludes that $\operatorname{var}\left(\mathcal{D}_{n}\right)$ is greater than a constant times $\log n$ for large enough $n$. This corresponds to the $a=b=1$, $\xi=0$ case. Compared with the bound in Lemma 4.4, we know the order of $c_{n}$ cannot be improved in general: We required $c_{n}^{2}$ to be slightly larger than $\log n$, and the latter matches the lower bound of ${\rm var}(\mathcal{D}_{n})$.
@@
This corresponds to the $a=b=1$, $\xi=2$ case, and we can compare the bound in Lemma 4.4.
In Figure D.1, we present the histograms of $\mathcal{D}_{n}$ for the mean estimation example with $\xi=1$ and $2$. The variances can be shown to be larger than $n^{2}$ and $n^{4}$, respectively, using the argument above. After dividing them by their standard deviation (SD)—more accurately, a sequence of the same order as their SD lower bounds—we observe their normalized distributions do not visually vary between different sample sizes (and are close to a tamed continuous distribution). Given the above evidence, we conjecture $c_{n}$ in Lemma 4.4 needs to be of order larger than $\sqrt{{\rm var}(\mathcal{D}_{n})}$ to have the probability in (16) converging to zero.
@@
$\mathbb{E}\left[\nabla_{j}u_{i+1}\mid F^{j}\right]=i^{\xi-2}S_{j}$
@@
$S_{j}=2\left(Y_{j}-\mu\right)\left(\sum_{l=1}^{j-1}Y_{l}-j\mu\right)+\left(Y_{j}^{2}-\alpha^{2}\right).$
@@
$\mathbb{E}\left[\nabla_{j}u_{i+1}\mid F^{j}\right]$
$=i^{\xi-2}\mathbb{E}\left[\left(Y_{j}-Y_{j}^{\prime}\right)\left(2\sum_{l=1,l\neq j}^{i}Y_{l}+Y_{j}+Y_{j}^{\prime}-2i\mu\right)\mid F^{j}\right]$
$=i^{\xi-2}Y_{j}\mathbb{E}\left[2\sum_{l=1,l\neq j}^{i}Y_{l}+Y_{j}+Y_{j}^{\prime}-2i\mu\mid F^{j}\right]$
$\quad-i^{\xi-2}\mathbb{E}\left[Y_{j}^{\prime}\left(2\sum_{l=1,l\neq j}^{i}Y_{l}+Y_{j}+Y_{j}^{\prime}-2i\mu\right)\mid F^{j}\right]$
@@ Appendix E Consistency under Diverging Candidate Cardinality
We define $H_{n}=\sum_{i=1}i^{\xi}\min_{k\in K_{i}}\mathbb{E}[r(h_{i-1}^{(k)})]$ as the accumulated error of the best alternative models.
( correct selection ) $\mathbb{P}(\text{ correct selection })$ (40)
$=$ $\mathbb{P}\left(\inf_{k\in K_{n}}\sum_{i=1}^{n}i^{\xi}\left\{h^{(k)}_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\right)$
$\geq$ $\mathbb{P}\left(\inf_{k}\sum_{i=1}^{n}i^{\xi}\left\{h^{(k)}_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\text{ and }\right.$
@@
$=$ $1-\mathbb{P}\left(\inf_{k}\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right.\text{ or }$
@@
$\geq$ $1-\mathbb{P}\left(\inf_{k}\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$
$\quad-\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$
$\geq$ $1-\sum_{k\in[K_{n}]}\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$
$\quad-\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right).$
Define $H_{n}^{(k)}=\sum_{i=1}^{n}i^{\xi}\mathbb{E}[r(h_{i-1}^{(k)})]$, for each term in the $\sum_{k\in[K_{n}]}$ summation:
$\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)$ (41)
$\leq\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}^{(k)}\right)$
$\leq\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}^{(k)}\right|\geq\delta H_{n}^{(k)}\right)$
$\stackrel{{\scriptstyle(I)}}{{\leq}}\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}^{(k)}\right|\geq\delta\sum_{i=1}^{n}i^{-a_{k}+\xi}M_{k}\right)$
$\leq\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\left\{h^{(k)}_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}-H_{n}^{(k)}\right|\geq\delta n^{1-a_{k}+\xi}\tilde{M}_{k}/2\right)$
@@
$\stackrel{{\scriptstyle(II)}}{{\leq}}\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)+$
$\quad\mathbb{P}\left(\left|\sum_{i=1}^{n}u_{i}^{(k)}-\mathbb{E}\left[u_{i}^{(k)}\mid F^{i-1}\right]\right|\geq c_{n}\right)+$
$\quad\mathbb{P}\left(\left|\sum_{i=1}^{n}\mathbb{E}\left[u_{i}^{(k)}\mid F^{i-1}\right]-H_{n}^{(k)}\right|\geq c_{n}\right).$
@@
$\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)\lesssim\exp(-K^{\prime}(\mathcal{C}^{(k)}_{n})^{\theta^{\prime}}),$ (42)
@@
$\mathcal{C}^{(k)}_{n}=n^{(1-a_{k})/2}\wedge\log^{-1}n\left(n^{b_{k}-a_{k}/2-1/2}\wedge n^{1-a_{k}+\xi}\right).$
@@
$\log^{-1}n\cdot n^{\mathcal{C}_{n}/2}\lesssim\mathcal{C}^{(k)}_{n}.$